# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v2

@@ 5 Conclusions and prospects
In this paper, we improve the PAC method originally developed in Paper I by introducing several enhancements to make it more rigorous and to eliminate the need for redshift bins.
@@
As a first application, we use the DECaLS photometric sample and the BGS Bright spectroscopic sample from DESI Y1 to measure $\bar{n}_{2}w_{{\rm{p}}}$ across various stellar mass bins, ranging from $10^{5.3}{\rm M}_{\sun}$ to $10^{11.5}{\rm M}_{\sun}$ for blue galaxies and $10^{6.3}{\rm M}_{\sun}$ to $10^{11.9}{\rm M}_{\sun}$ for red galaxies, all of which are complete in stellar mass. Additionally, we measure several incomplete bins down to $10^{3.7}{\rm M}_{\sun}$ for blue galaxies and $10^{5.5}{\rm M}_{\sun}$ for red galaxies. Following Paper III with some extensions, we combine these $\bar{n}_{2}w_{{\rm{p}}}$ results with the corresponding $w_{{\rm{p}}}$ measurements from the BGS sample, which may not be fully complete in stellar mass. Assuming that galaxy bias is primarily determined by stellar mass and colour, we then calculate $\bar{n}_{2}$ for these mass bins. This approach enables us to derive the GSMF down to $10^{5.3}{\rm M}_{\sun}$ for blue galaxies and $10^{6.3}{\rm M}_{\sun}$ for red galaxies, with lower limits extending down to $10^{3.7}{\rm M}_{\sun}$ and $10^{5.5}{\rm M}_{\sun}$, respectively. We also examine the influence of fragmented sources in the DECaLS samples on our GSMF measurements, which could be misidentified features like fragmented spiral arms or star-forming regions. The impact on red galaxies is negligible. However, for blue galaxies, the fake fraction $F_{\rm{frag}}$ is notably higher at $M_{*}<10^{9.0}{\rm M}_{\sun}$, though it remains below $5\%$ for $M_{*}>10^{5.0}{\rm M}_{\sun}$. For $M_{*}<10^{5.0}{\rm M}_{\sun}$, $F_{\rm{frag}}$ becomes significant, reaching $40\%$ at $10^{4.6}{\rm M}_{\sun}$. These fake sources have been corrected for in our GSMF.
@@
We find that the GSMFs for blue, red, and the entire galaxy populations can be well described by single, double, and triple Schechter functions. The power-law indices at the low-mass ends are $-1.54^{+0.02}_{-0.02}$ for blue galaxies and $-2.50^{+0.02}_{-0.02}$ for red galaxies, resulting in an index ranging between $-1.85$ and $-2.56$ for the entire GSMF. The slope of our GSMF at the low-mass end is significantly steeper than that of GAMA (Driver et al., 2022) and BGS Bright , whose index is closer to that of our blue GSMF. This discrepancy is reconciled after applying the correction for the local volume, as provided by Chen et al. (2019), to the BGS $V_{\rm{max}}$ GSMF. Additionally, the steep rise in the GSMF at lower masses has not been successfully reproduced by current galaxy formation models, suggesting an incomplete understanding of the physics governing galaxy formation.
@@
Our measurements of the GSMF, extending to the $10^{6.0}{\rm M}_{\sun}$ range, hold the potential to provide valuable constraints on various aspects, including the nature of dark matter (Colín et al., 2000; Bode et al., 2001; Viel et al., 2005; Lovell et al., 2012; Macciò et al., 2013; Horiuchi et al., 2016; Bose et al., 2017), the epoch of reionization (Efstathiou, 1992; Loeb & Barkana, 2001; Okamoto et al., 2008; Benítez-Llambay et al., 2017; Bose et al., 2018), and the physics of galaxy formation. For instance, warm dark matter models predict a cutoff at high $k$ in the power spectrum, leading to the suppression of both the halo and stellar mass functions at low masses (Bose et al., 2017). The steep rise we observe in our GSMF at the low-mass end may place strong constraints on the mass of dark matter particles. Furthermore, upcoming cosmological photometric surveys, such as LSST, Euclid, CSS-OS, and Roman, are expected to probe depths 2-3 magnitudes fainter than DECaLS. By applying the PAC method, we have the potential to thoroughly explore the full range of the galaxy population in the local universe and address these fundamental questions.
@@
In addition to its relevance for dwarf galaxy studies in the local universe, the PAC method offers many other applications. By leveraging the diverse spectroscopic samples from DESI, including BGS, LRG, ELG, and QSO, which together map the universe up to $z\sim 2$, we can consistently investigate the evolution of various galaxy properties. Moreover, we can establish highly accurate galaxy-halo connections for spectroscopic tracers, thanks to the extensive information available on their spatial distribution in relation to different types of galaxies, as developed in Paper IV. This enables precise modeling of small-scale measurements based on these tracers, such as galaxy-galaxy lensing and redshift space distortions (Li et al., 2012; Yuan et al., 2022; Zhai et al., 2023; Xu et al., 2024), leading to stringent constraints on the structure formation of the universe.

### v3

@@ 5 Conclusions and prospects
In this paper, we improve the PAC method originally developed in 2022ApJ...925...31X by introducing several enhancements to make it more rigorous and to eliminate the need for redshift bins.
@@
As a first application, we use the DECaLS photometric sample and the BGS Bright spectroscopic sample from DESI Y1 to measure $\bar{n}_{2}w_{{\rm{p}}}$ across various stellar mass bins, ranging from $10^{5.3}{\rm M}_{\sun}$ to $10^{11.5}{\rm M}_{\sun}$ for blue galaxies and $10^{6.3}{\rm M}_{\sun}$ to $10^{11.9}{\rm M}_{\sun}$ for red galaxies, all of which are complete in stellar mass. Additionally, we measure several incomplete bins down to $10^{3.7}{\rm M}_{\sun}$ for blue galaxies and $10^{5.5}{\rm M}_{\sun}$ for red galaxies. Following 2022ApJ...939..104X with some extensions, we combine these $\bar{n}_{2}w_{{\rm{p}}}$ results with the corresponding $w_{{\rm{p}}}$ measurements from the BGS sample, which may not be fully complete in stellar mass. Assuming that galaxy bias is primarily determined by stellar mass and colour, we then calculate $\bar{n}_{2}$ for these mass bins. This approach enables us to derive the GSMF down to $10^{5.3}{\rm M}_{\sun}$ for blue galaxies and $10^{6.3}{\rm M}_{\sun}$ for red galaxies, with lower limits extending down to $10^{3.7}{\rm M}_{\sun}$ and $10^{5.5}{\rm M}_{\sun}$, respectively. We also examine the influence of fragmented sources in the DECaLS samples on our GSMF measurements, which could be misidentified features like fragmented spiral arms or star-forming regions. The impact on red galaxies is negligible. However, for blue galaxies, the fake fraction $F_{\rm{frag}}$ is notably higher at $M_{*}<10^{9.0}{\rm M}_{\sun}$, though it remains below $5\%$ for $M_{*}>10^{5.0}{\rm M}_{\sun}$. For $M_{*}<10^{5.0}{\rm M}_{\sun}$, $F_{\rm{frag}}$ becomes significant, reaching $40\%$ at $10^{4.6}{\rm M}_{\sun}$. These fake sources have been corrected for in our GSMF.
@@
We find that the GSMFs for blue, red, and the entire galaxy populations can be well described by single, double, and triple Schechter functions. The power-law indices at the low-mass ends are $1.54^{+0.02}_{-0.02}$ for blue galaxies and $2.50^{+0.02}_{-0.02}$ for red galaxies, resulting in an index ranging between 1.85-1.85 and 2.56-2.56 for the entire GSMF. The slope of our GSMF at the low-mass end is significantly steeper than that of GAMA (2022MNRAS.513..439D) and BGS Bright , whose index is closer to that of our blue GSMF. This discrepancy is reconciled after applying the correction for the local volume, as provided by 2019ApJ...872..180C, to the BGS $V_{\rm{max}}$ GSMF. Additionally, the steep rise in the GSMF at lower masses has not been successfully reproduced by current galaxy formation models, suggesting an incomplete understanding of the physics governing galaxy formation.
@@
Our measurements of the GSMF, extending to the $10^{6.0}{\rm M}_{\sun}$ range, hold the potential to provide valuable constraints on various aspects, including the nature of dark matter (2000ApJ...542..622C; 2001ApJ...556...93B; 2005PhRvD..71f3534V; 2012MNRAS.420.2318L; 2013MNRAS.428..882M; 2016MNRAS.456.4346H; 2017MNRAS.464.4520B), the epoch of reionization (1992MNRAS.256P..43E; 2001ARA&A..39...19L; 2008MNRAS.390..920O; 2017MNRAS.465.3913B; 2018ApJ...863..123B), and the physics of galaxy formation. For instance, warm dark matter models predict a cutoff at high $k$ in the power spectrum, leading to the suppression of both the halo and stellar mass functions at low masses (2017MNRAS.464.4520B). The steep rise we observe in our GSMF at the low-mass end may place strong constraints on the mass of dark matter particles. Furthermore, upcoming cosmological photometric surveys, such as LSST, Euclid, CSS-OS, and Roman, are expected to probe depths 2-3 magnitudes fainter than DECaLS. By applying the PAC method, we have the potential to thoroughly explore the full range of the galaxy population in the local universe and address these fundamental questions.
@@
In addition to its relevance for dwarf galaxy studies in the local universe, the PAC method offers many other applications. By leveraging the diverse spectroscopic samples from DESI, including BGS, LRG, ELG, and QSO, which together map the universe up to $z\sim 2$, we can consistently investigate the evolution of various galaxy properties. Moreover, we can establish highly accurate galaxy-halo connections for spectroscopic tracers, thanks to the extensive information available on their spatial distribution in relation to different types of galaxies, as developed in 2023ApJ...944..200X. This enables precise modeling of small-scale measurements based on these tracers, such as galaxy-galaxy lensing and redshift space distortions (2012ApJ...758...50L; 2022MNRAS.515..871Y; 2023ApJ...948...99Z; 2024ApJ...973..102X), leading to stringent constraints on the structure formation of the universe.